OpenStudy (anonymous):

Hey can you please answer this? Each time you click a toggle switch, the switch either turns from off to on or from on to off . Suppose that you start with three toggle switches with one of them on and two of them off . On each move you randomly select one of the three switches and click it. Let m and n be relatively prime positive integers so that m/n is the probability that after four such clicks, one switch will be on and two of them will be off . Find m + n.

OpenStudy (anonymous):

i don't think this will be easy, and i think it will take some time

OpenStudy (anonymous):

did you try anything? for example, if all the moves are on one switch, then there will be one on and two off right?

OpenStudy (anonymous):

likewise if three are on the first switch (the on one) and the other one is not , then there are still two off and one on

OpenStudy (anonymous):

if two are on the first switch, then the next two must be on one, otherwise you get on on on

OpenStudy (anonymous):

after going through some cases, maybe it would be easier to find the probability that there is one off and two on. seem less likely

OpenStudy (anonymous):

meant the other way around, sorry one off, two on

OpenStudy (anonymous):

i found out that its impossible that after 1 click and 3 clicks, it is impossible to have the same combination as the first time

OpenStudy (anonymous):

if all four on one switch, one on and two off if three on the on switch, then still one off and two on if two from the on switch, then you have on off off and so the next two have to be from different off switches that probability is \((\frac{1}{3})^4\) two from the first, one each from the second

OpenStudy (anonymous):

also found that after 2 clicks, it is 1/3 probability to have them good

OpenStudy (anonymous):

to summerize so far, all four from any one switch, one on and two off three from first , one on and two off two from first only get one off and two on if they are from different switches \(p=(\frac{1}{3})^4\) what about one from the first switch?

OpenStudy (anonymous):

one on the first switch, then the other two have to come from different switches to get all off

OpenStudy (anonymous):

probability of that is also \(p=(\frac{1}{3})^4\)

OpenStudy (anonymous):

so far only two ways not to end up with one on and two off, both with probability \((\frac{1}{3})^4\)

OpenStudy (anonymous):

how about none on the first switch?

OpenStudy (anonymous):

then must be three on one and one on the other, othewise still one on and two off

OpenStudy (anonymous):

probability of that is \(2\times (\frac{1}{3})^4\)

OpenStudy (anonymous):

i think that is it unless i miscounted

OpenStudy (anonymous):

so although i would not bet on it, looks like total ways to get them not one on and two off is \[4\times (\frac{1}{3})^4=\frac{4}{81}\] and so probability of one on and two off is \[\frac{77}{81}\]

OpenStudy (anonymous):

this may certainly be wrong, but if you check through the cases i think that is the method needed

OpenStudy (anonymous):

im currently checking that out, thanks for you help :)